OpenStudy (anonymous):

solve..integration sign--> S (2+ln u )/u^2 by substitution u = e^x

OpenStudy (amistre64):

this one?

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

since u = e^x; what does u' equal?

OpenStudy (anonymous):

e^x i guess

OpenStudy (amistre64):

yep, so we need to substitute in all the proper parts \[\int \frac{2+ln(u)}{u^2}du\] \[u=e^x\]\[du = e^x~dx\] \[\int \frac{2+ln(e^x)}{(e^x)^2}e^x~dx\] can you simplify that ?

OpenStudy (anonymous):

how did you get e^x dx

OpenStudy (amistre64):

it is given that u=e^x ; to find out what du equals, we have to take the derivative of u

OpenStudy (anonymous):

thnx i got it i was jst missing e^x dx part therefore getting wrong answer may thnx

OpenStudy (anonymous):

many thnx

OpenStudy (amistre64):

\[\frac d{dx}(u=e^x)\] \[\frac d{dx}(u)=\frac d{dx}(e^x)\] \[\frac {du}{dx}=e^x\] \[du=e^x~dx\]

OpenStudy (amistre64):

yeah, you gotta make sure you are replacing all the u parts with their correct x counter parts ;) good luck